OpenStudy (anonymous):

triangle ABC, the measure of angle A is 24degrees more than the measure of angle B. The measure of angle B is the same as the measure of angle ACB. Find the measures of the 3 interior angles and the exterior angle ACD

OpenStudy (anonymous):

wouldnt <A be 24 +x

OpenStudy (yttrium):

yeah, i mean 24+X

OpenStudy (yttrium):

Let <B =x = <ACB <A = 24 + x Therefore, <A+<B+<C=180 (24+x)+x+x = 180 Solve for x and then substitute the value of it to each interior angle. For exterior angle ACD: <ACD = 360 - <ACB

OpenStudy (anonymous):

so x=52 ?

OpenStudy (yttrium):

yes.:)

OpenStudy (goformit100):

First of all @Cardmich93 and @Yttrium : A Warm Welcome to 'Open Study'.

OpenStudy (anonymous):

And <ACD is 308

OpenStudy (yttrium):

Yes you got it right! :D

OpenStudy (anonymous):

ok thanks so much, now 5 more questions to go :)

OpenStudy (yttrium):

that's fine. :))